Part A Item 10 What is the area of the loop? Express your answer using two significant figures A uniform magnetic field is at right angles to the plane of a wire loop. The field decreases by 0.20 T in 3.0x10s and the magnitude of the average emf induced in the loop is 77 V Submit My Answers Give Up Part B What would be the value of the average induced emf if the field change was the same but took twice as long to decrease? Express your answer using two significant figures Submit My Answers Give Up Part C What would be the value of the average induced emf if the field decrease was twice as much and it also took twice as long to change? Express your answer using two significant figures
Answer #1

A) induced emf = rate of change of flux

        e = d(NBA)dt = NAdB/dt

         77 = 1* A * (0.20)/(3 x 10^-3)

       A = 1.155 m^2

b) now dt is doubled then e wil halved.

e = 77/2 = 38.5 Volt

c) then dB/dt willl not change.

so e = 77 volt